OpenStudy (anonymous):

The radius r of a sphere is increasing at a rate of 9 inches per minute. (a) Find the rate of change of the volume when r = 10 inches. .......... in.3/min (b) Find the rate of change of the volume when r = 36 inches. .............. in.3/min

OpenStudy (anonymous):

But this is not the right answer for some reason can someone help me?

OpenStudy (astrophysics):

Related rates, so what did you try?

OpenStudy (astrophysics):

Volume of a sphere is \[V = \frac{ 4 }{ 3 } \pi r^3\] and we know \[\frac{ dr }{ dt } = 9 \frac{ i n }{ \min }\]

OpenStudy (astrophysics):

We need to find dV/dt

OpenStudy (anonymous):

yes i got to the point where dV/dt = 4 pi r ^2 (dR/dt)

OpenStudy (anonymous):

and then put in the numbers and got 3600 but it isn't right

OpenStudy (astrophysics):

\[\frac{ dV }{ dt } = 4 \pi r^2 \frac{ dr }{ dt } \implies 4 \pi (10i n)^2(9) = 3600 \pi \frac{ i n^3 }{ \min }\]

OpenStudy (anonymous):

oh lol thanks

OpenStudy (astrophysics):

forget pi? :P

OpenStudy (anonymous):

yeah

OpenStudy (astrophysics):

:)

OpenStudy (astrophysics):

Same process with r = 36

OpenStudy (anonymous):

yep thanks